Let P(E)= 0.37, P(EF)= 0.19, and P(EF^c)= 0.89. Find P(F|E^C).
Answer #1

P(E)=0.37

PE=1-PE) = 1-0.37 = 0.63

P(E)+P(F)=1

P(F)=1-0.37=0.63

PENF) = 0.19

PENF) = 0.89

P(F|EC) - P(FNE) P(EC)

P(FO E) = P(F) - P(FIE)

P( F E ) = 0.63 - 0.19 = 0.44

P(FE) = P(F n E“) P(E) 0.44 062 = 0.698